OpenStudy (anonymous):

Solve the equation -2(5 + 6c) + 16 = -90

OpenStudy (anonymous):

-10 + -12c = -90 -12c= -80 12c=80 c = 6.666666667

OpenStudy (anonymous):

Thanks

Parth (parthkohli):

= (-10 - 12c) + 16 = -90 = -10 - 12c = -106 = -12c = -96 = 12c = 96 = c = 8

Parth (parthkohli):

c is 8 not 6.66666666667

OpenStudy (anonymous):

Okay , lol